Fermi National Laboratory

Accelerator Update

Friday October 8


    Bo Low Beta Separators HLS Installation - 10%

    A0 Vacuum Upgrade - 60%

    Conning Tower work at E2, F2, & F3 - 100%

    E4 Vacuum Work - 100%

    IPM and Crystal Collimator installation - 50%

    D4 Spool Replacement - 100%

    D17 Separator Installation (discovered a vacuum leak) - 75% (Old Info)

    D0 Low Beta Separators HLS Installation - 10%

    Shimming 328 Dipoles - 65% complete (Old Info)


    Drift Tube Tank #5 Replacement - 100%, should be under vacuum 10/8/04


    RF19 - 90% complete, should be done around 10/13/04

    Long 13 Work - 100% complete (Long 13 is a sector of the Booster ring.)

Antiproton Source

    Debuncher Injection Area Beampipe and Quadrupole (Quad) replacement - Checkout in progress. Work on the last four D40 quads are in progress. Septum magnet pulse tested at full current. The beam pipe is almost ready. The kicker girder in progress

      Debuncher Motorized Quads Installation - ~80% complete

    Accumulator Flying Wires - The wires have been installed and tested. They are waiting on vacuum pump down.

    ARF1-1 Installation - ~95% complete. Alignment completed.

    AP40 - Replaced vertical flying wire. Bake out will begin on Monday

    AP10 SO Cord Remediation - 75% complete. AP10, F23, and F27 work in progress.

    D/A Lines -alignment check in progress, cooling system measurements next week, visual inspections after "as founds", and septum pulse testing after "as founds"

    Beam Lines - Work in progress: AP1 magnet moves, new VT108, reconfigure vacuum windows, AP2 upstream vacuum will be opened next week.

    Other Work - Target Station, Alignment, Stochastic Cooling, Controls, Cryo, LCW, and LLRF - all work on track

Main Injector/Recycler
    Overall shut down progress - 70% (Old Info)

    MI LCW system is up and running.


      Q307-309 - 7% complete (Old Info)

    MI-30 & 62


      Cooling Tank Repairs - ~75% complete (Old Info)
    Horn Replacement - 61%, this work is one week behind due to many problems
    The kicker is installed and ready. The Recycler shielding is near to completion. The trim magnets are all in place. Alignment is half done and about one week behind schedule due to equipment trouble.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/11/2004   email Fermilab